OpenStudy (anonymous):

A building lot in a city is shaped as a 30° -60° -90° triangle. The side opposite the 30° angle measures 41 feet. a. Find the length of the side of the lot opposite the 60° angle. b. Find the length of the hypotenuse of the triangular lot. c. Find the sine, cosine, and tangent of the 30° angle in the lot. Write your answers as decimals rounded to four decimal places.

OpenStudy (anonymous):

please helppp

OpenStudy (anonymous):

1. tan(30) = 41/x --> x = 41/tan(30) where x is the side opposite the 60 degree angle x = 71.0141 ft 2.Tangent (30) = 41/a a = 41/0.5773 = 71 feet \[Hypotenuse = \sqrt{a^2+b^2}\] \[\sqrt{5041 + 1681}=82\] 3. sine 30° = 1/2 = 0.5000 cosine 30° = \[\sqrt{3/2}= 0.8660\] tangent 30° = \[1/\sqrt{3}\]= 0.5774 @DianaNicolee01